# Topology and Algebra of Bonded Knots and Braids

### v2

@@ 4.2. Rigid Standard Bonded Links: Tangle Insertions
  • Into each band, insert a properly embedded 222–tangle chosen from a fixed family of tangles.
@@
[…]
@@ 4.3. Rigid Tight Bonded Links: The Bonded Bracket Polynomial
In this subsection we present the bonded bracket polynomial, which extends the Kauffman bracket polynomial to rigid tight bonded links by assigning specific skein-theoretic relations at bonds. This invariant arises by choosing a particular 222–tangle insertion at each bond, corresponding to a fixed linear combination of possible 222–tangle states, weighted by formal coefficients. The resulting classical diagram (obtained after resolving all bonds via this prescribed 222–tangle insertion) is then evaluated using the usual Kauffman bracket polynomial.
We emphasize that the construction described below uses one specific choice of 222–tangle insertion at each bond. In principle, different choices of 222–tangle insertions give rise to different invariants. Indeed, one can think of this invariant as a representative example of a much broader class of invariants defined by varying the 222–tangle substitutions at bonds. Exploring how powerful this family of invariants is remains an open problem.
@@
The bonded bracket polynomial is an invariant of regular isotopy in the category of tight rigid bonded links.
@@
Figure 39. Invariance under a node R2-type move.
@@
To illustrate the bonded bracket polynomial, we explicitly compute it for two families of examples, corresponding to specific choices of inserted tangles at the bond, and, for simplicity, we set $b=1$.
@@
The proof follows by induction, as illustrated in Figures 40 and 41, taking into account that the addition of a local curl to a bonded diagram by an R1 move multiplies the bracket polynomial of the diagram by a factor of $-A^{3}$ if the curl is right-handed and $-A^{-3}$ if the curl is left-handed.
@@
We prove relations (1) by induction on $n\in\mathbb{N}\backslash\{0\}$. For $n=1$ we have that:
@@
We note that the highest power of $a$ appearing among all monomials in the expansion of the bonded bracket polynomial of a bonded link diagram equals the total number of bonds.
Finally, we note that invariants defined for standard bonded links extend to tight bonded links, and, in some cases, they become simpler to compute due to the absence of crossings between bonds and link arcs.
@@ 5. Bonded Braids
We recall that a geometric braid on $n$ strands is a homeomorphic image of $n$ arcs in the interior of $[0,1]\times[0,\epsilon]\times[0,1]$, $\epsilon>0$, such that it is monotonous, that is, there are no local maxima or minima, and the boundary of the image consists in $n$ numbered points in $[0,1]\times[0,\epsilon]\times\{0\}$ and $n$ corresponding numbered points in $[0,1]\times[0,\epsilon]\times\{1\}$. We study braids through their diagrams in the plane $\}\times[0,1]$ 0 } × [ 0 , 1 ], which are also called braids. Formally, a braid is described by its braid word in the braid group $B_{n}$, which has generators $\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1}$ and relations:
@@ 5.1. Bonded braid definition
A bonded braid on $n$ strands is a pair $(\beta,B)$ of a classical braid $\beta$ on $n$ strands, and a set of $k$ disjoint, embedded horizontal simple arcs, called bonds. The boundary points of a bond, called nodes, cannot coincide with endpoints of the braid or with other nodes, and they have local neighborhoods that are 3-valent graphs with the attaching arcs, like $\vdash$, forming together an H-neighborhood, as depicted in Figure 23. A bond joining the $i^{th}$ and $j^{th}$ strands with $i<j$ is denoted $b_{i,j}$ and it threads transversely through the strands in between. By abuse of notation, we denote by $b_{i,j}$ a bond joining the $i^{th}$ and $j^{th}$ strands with any sequence of overpasses or underpasses. A bond $b_{i,i+1}$ joining two consecutive strands $i$ and $i+1$ shall be called elementary bond, and will be denoted by $b_{i}$. If $B=\emptyset$, then $(b,\emptyset)$ is just a classical braid. Moreover, a bonded braid diagram (also referred to as bonded braid) is a regular projection of a bonded braid in the plane $\}\times[0,1]$ 0 } × [ 0 , 1 ] with over/under information at every crossing, and such that no crossing is horizontally aligned with a bond. So, $b_{i,j}$ can be encoded unambiguously by a sequence of $j-i+1$ o’s and u’s indicating the types of crossings formed between the bond and any strand in between. For an example see left-hand side illustration of Figure 44. A configuration of a bond which passes either over or under all its intermediate strands, such that all crossings are marked ‘over’/‘under’, is a uniform over/under.
@@ 5.2. Bonded braid isotopy
    Figure 46. Interactions between two bonds that are far away from each other.
@@
[…]
@@
A interaction between a bond and a crossing depends on the relative positions of the bond and the strands forming the crossing. More precisely:
@@
    Figure 50. A crossing not involving a bond.
@@
  • Figures 53, 54 illustrate the cases where one node of the bond lies on a strand, resp. on two stands, that crosses (resp. cross) another strand. In the instance of Figure 54 there may be other strands threading through the bond, as long as they all lie under the overcrossing strand.
@@
The additional moves depicted in Figures 47, 48 and 52 are essential to define bonded braid equivalence. These moves account for the interactions specific to bonds, ensuring that the topological properties of the bonded braid are preserved. Using these moves, we extend classical braid isotopy to include bonded braid isotopy, leading to the following formal definition.
@@ 5.3. The Bonded Braid Monoid
$\sigma_{1},\ldots,\sigma_{n-1}\quad(\text{classical braid generators})\quad\text{and}\quad b_{i,j}\quad(1\leq i<j\leq n)\quad(\text{bonds}),$
@@
$\begin{array}[]{rcll}\sigma_{i}\,\sigma_{j}&=&\sigma_{j}\,\sigma_{i}&\quad\text{for }\ |i-j|>1,\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\quad\text{for all }\ i,\\[4.0pt] b_{i}\,b_{j}&=&b_{j}\,b_{i}&\quad\text{for }|i-j|>1,\\[4.0pt] b_{i}\,{\sigma_{j}}^{\pm 1}&=&{\sigma_{j}}^{\pm 1}\,b_{i}&\quad\text{for }\ |i-j|>1,\\[4.0pt] b_{i}\,\ {\sigma_{i}}^{\pm 1}&=&{\sigma_{i}}^{\pm 1}\,b_{i}&\quad\text{for all }\ i,\\[4.0pt] b_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,b_{i+1}&\quad\text{for all }\ i,\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,b_{i}&=&b_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\quad\text{for all }\ i.\end{array}$
@@
For reference, the singular braid monoid $SB_{n}$ is generated by the standard braid generators $\sigma_{1}^{\pm 1},\ldots,\sigma_{n-1}^{\pm 1}$ together with singular crossing generators $\tau_{1},\ldots,\tau_{n-1}$, subject to the following relations:
$\begin{array}[]{rcll}\sigma_{i}\,\sigma_{j}&=&\sigma_{j}\,\sigma_{i}&\text{for }|i-j|>1,\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\text{for all }i,\\[4.0pt] \tau_{i}\,\tau_{j}&=&\tau_{j}\,\tau_{i}&\text{for }|i-j|>1,\\[4.0pt] \tau_{i}\,{\sigma_{j}}^{\pm 1}&=&{\sigma_{j}}^{\pm 1}\,\tau_{i}&\text{for }|i-j|>1,\\[4.0pt] \tau_{i}\,{\sigma_{i}}^{\pm 1}&=&{\sigma_{i}}^{\pm 1}\,\tau_{i}&\text{for all }i,\\[4.0pt] \tau_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,\tau_{i+1}&\text{for all }i\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,\tau_{i}&=&\tau_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\text{for all }\ i.\\[4.0pt] \end{array}$
@@
$\begin{array}[]{rcl}b_{2}&=&(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})\\[4.0pt] b_{3}&=&(\sigma_{3}\sigma_{2})^{-1}(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})(\sigma_{3}\sigma_{2})\\[4.0pt] &\vdots&\\[4.0pt] b_{i}&=&(\sigma_{i}\sigma_{i-1})^{-1}\cdots(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})\cdots(\sigma_{i}\sigma_{i-1}),\\[4.0pt] &\vdots&\end{array}$
@@
The tight bonded braid monoid $BB_{n}$ admits the following irredundant presentation: it is generated by the classical braid generators $\sigma_{1},\ldots,\sigma_{n-1}$ and a single bond generator $b_{1}$, subject to the relations:
$\begin{array}[]{rcll}\sigma_{i}\,\sigma_{j}&=&\sigma_{j}\,\sigma_{i}&\quad\text{for }|i-j|>1,\\[4.0pt] \sigma_{i}\,\sigma_{i+1}\,\sigma_{i}&=&\sigma_{i+1}\,\sigma_{i}\,\sigma_{i+1}&\quad\text{for all }i,\\[4.0pt] b_{1}\,{\sigma_{j}}^{\pm 1}&=&{\sigma_{j}}^{\pm 1}\,b_{1}&\quad\text{for }j>2,\\[4.0pt] b_{1}\,(\sigma_{2}\sigma_{1})(\sigma_{1}\,\sigma_{2})&=&(\sigma_{2}\sigma_{1})(\sigma_{1}\,\sigma_{2})\,b_{1}&\\[4.0pt] b_{1}\,{\sigma_{1}}^{\pm 1}&=&{\sigma_{1}}^{\pm 1}\,b_{1}.&\end{array}$
6. The analogue of the Alexander Theorem for bonded links
In this section we focus on the theory of topological standard bonded links and we present a braiding algorithm for oriented standard bonded links.
Namely, we prove the following theorem:
@@ 6. The analogue of the Alexander Theorem for bonded links
Every oriented topological standard bonded link can be represented isotopically as the closure of a standard resp. tight bonded braid.
@@
Figure 59. Up-arcs.
@@
Step 2: We next deal with up-arcs in the diagram that contain at least one node of a bond. For this we apply isotopy (TVT or RVT) as shown in Figures 62 and 63, according to whether the two joining arcs are antiparallel or parallel upward arcs. Then the bonds will only lie on down-arcs, so that the braiding algorithm will not affect them.
Figure 62. Braiding two anti-parallel bonded strands.
Figure 63. Braiding two parallel bonded strands oriented upward.
@@ 7. Bonded braid equivalences
In this section we formulate and prove two bonded braid equivalences for topological bonded braids: the bonded $L$-equivalence and the analogue of the Markov theorem. For proving a braid equivalence we need to have the diagrams in general position, as described in the proof of the bonded braiding algorithm, to which we now add the triangle condition, whereby any two sliding triangles of up-arcs do not intersect. This is achieved by further subdivision of up-arcs, if needed, and it allows any order in the sequence of the braiding moves. Note that the interior of a sliding triangle may intersect a bond.
We proceed with introducing the notion of $L$-moves in the bonded setting. $L$-moves make up an important tool for formulating braid equivalences in practically any topological setting and they prove to be particularly useful in settings where the braid analogues do not have a ‘nice’ algebraic structure. In [33, 32, 25] $L$-moves and braid equivalence theorems are presented for different knot theories.
@@
An $L$-move on a bonded braid $\beta$, consists in the following operation or its inverse: we cut an arc of $\beta$ open and we pull the upper cutpoint downward and the lower upward, so as to create a new pair of braid strands with corresponding endpoints (on the vertical line of the cutpoint), and such that both strands run entirely over or entirely under the rest of the braid (including the bonds). Pulling the new strands over will give rise to an $L_{o}$-move and pulling under to an $L_{u}$-move. By definition, an $L$-move cannot occur in a bond or at a node. Figure 64 shows $L_{o}$ and $L_{u}$ moves taking place above or below a bond, on its attaching strand.
@@
Figure 65. A crossing formed after an $L$-move is performed, and resolving the $L$-move by pulling strands away from the bond.
@@
Theorem 7.3 ($L$-equivalence for topological bonded braids).
Two oriented topological standard bonded links are topologically isotopic if and only if for any two corresponding bonded braids one can be obtained from the other by a finite sequence of the following moves and bonded braid isotopy:
$\begin{array}[]{lllcll}1.&L-{moves}&&&&\\ 2.&{Bond\ Commuting:}&\alpha\,b_{i,j}&\sim&b_{i,j}\,\alpha,&{for}\ \alpha,\,b_{i,j}\in BB_{n}\\ \end{array}$
Furthermore, two oriented topological tight bonded links are topologically isotopic if and only if for any two corresponding tight bonded braids one can be obtained from the other by a finite sequence of the following moves and tight bonded braid isotopy:
$\begin{array}[]{lllcll}1.&{Tight\ L-moves}&&&&\\ 2.&{Elementary\ Bond\ Commuting:}&\alpha\,b_{i}&\sim&b_{i}\,\alpha,&{for}\ \alpha,\,b_{i}\in BB_{n}\\ \end{array}$
@@
Figure 66. Bond Commuting.
@@
For the one direction, the closures of two standard/tight bonded braids that differ by the moves (1) and (2) of either statement are clearly topologically vertex equivalent. For the converse, we need to show that any choices made for bringing a bonded diagram to general position result in $L$-equivalent bonded braids, and that any bonded isotopy moves on a bonded diagram correspond to $L$-equivalent bonded braids.
@@
We finally check the vertex slide (VS) moves. The three representing cases of oriented VS moves with the middle arc being an up-arc are illustrated in Figures 71, 72 and 73. The moves are considered local so that all other braiding is done and we can compare the final braids. Figure 71 shows the case of parallel down-arcs at the bonding sites. After the performance of the $L$-moves we see that the move is invisible on the bonded braid level.
@@
The $L$-moves naturally generalize the stabilization moves for classical braids, since an $L$–move is equivalent to adding a positive or a negative crossing (recall Note 7.2 and Figure 65). Moreover, as shown in [33], the $L$-moves can also realize conjugation for classical braids and this carries through also to bonded braids. On the other hand, as shown in Figure 65, an $L$–move can be created by braid isotopy, stabilization and conjugation. Hence, we may replace moves (1) of Theorem 7.3 and obtain the analogue of the Markov theorem for bonded braids.
Theorem 7.5 (Markov equivalence for bonded braids).
The closures of two bonded braids are topologically isotopic oriented standard bonded links if and only if one can be obtained from the other by a finite sequence of the following moves and braid isotopy:
$\begin{array}[]{lllcll}1.&{Markov\ Conjugation:}&\alpha&\sim&\sigma_{i}^{\pm 1}\,\alpha\,\sigma_{i}^{\mp 1},&{for}\ \alpha,\,\sigma_{i}\in BB_{n},\\ &&&&&\\ 2.&{Elementary\ Bond\ Commuting:}&\alpha\,b_{i}&\sim&b_{i}\,\alpha,&{for}\ \alpha,\,b_{i}\in BB_{n},\\ &&&&&\\ 3.&{Markov\ Stabilization:}&\alpha&\sim&\alpha\,\sigma_{n}^{\pm 1},&{for}\ \alpha\in BB_{n}.\\ \end{array}$
Note 7.6.
Note that a stabilization move involving bonds (see Figure 74) does not appear in the equivalence of Theorem 7.5. The reason is that such a ‘Markov bonded-stabilization move’ results in a vertical bond, which is out of the bonded braid category. The closure of the figure on the right can be converted by braiding moves to a legitimate bonded braid with a horizontal bond. Call this bonded braid diagram $P$. Then the left hand side and $P$ can be shown to be $L$-equivalent. We leave this as an exercise for the reader.
Figure 74. Markov stabilization moves involving bonds.

### v4

@@ 4.2. Rigid Standard Bonded Links: Tangle Insertions
  • Into each band, insert a properly embedded 22–tangle chosen from a fixed family of tangles.
@@
  • For any fixed choice of tangles, the ambient (resp. regular) isotopy class of the resulting classical link $L_{(K,B)}$ is a rigid vertex (resp. rigid vertex regular) isotopy invariant of $(K,B)$.
@@ 4.3. Rigid Tight Bonded Links: The Bonded Bracket Polynomial
In this subsection we present the bonded bracket polynomial, which extends the Kauffman bracket polynomial to rigid tight bonded links by assigning specific skein-theoretic relations at bonds. This invariant arises by choosing a particular 22–tangle insertion at each bond, corresponding to a fixed linear combination of possible 22–tangle states, weighted by formal coefficients. The resulting classical diagram (obtained after resolving all bonds via this prescribed 22–tangle insertion) is then evaluated using the usual Kauffman bracket polynomial.
We emphasize that the construction described below uses one specific choice of 22–tangle insertion at each bond. In principle, different choices of 22–tangle insertions give rise to different invariants. Indeed, one can think of this invariant as a representative example of a much broader class of invariants defined by varying the 22–tangle substitutions at bonds. Exploring how powerful this family of invariants is remains an open problem.
@@
The bonded bracket polynomial is an invariant of regular rigid vertex isotopy in the category of tight rigid bonded links.
@@
Figure 39. Invariance under a vertex slide move.
@@
To illustrate the bonded bracket polynomial, we explicitly compute it for three examples where, for simplicity, we set $b=1$.
@@
The proof follows by induction, as illustrated in Figures 40 and 41, taking into account that the addition of a local curl to a bonded diagram by an R1 move multiplies the bracket polynomial of the diagram by a factor of $A^{3}$ if the curl is right-handed and $A^{-3}$ if the curl is left-handed.
@@
[…]
@@
Example 4.13.
In this example we consider in Figure 43 two versions of bonded trefoil knots $T_{1}$ and $T_{2}$. In the figure the bonds are represented by curly black arcs.
Figure 43. Two bonded trefoils.
By expanding the crossings in these examples via the Kauffman bracket identity, we find that
$\langle T_{1}\rangle=(A^{3}d+3A^{2}B+AB^{2}d)\langle\Theta_{1}\rangle+(2AB^{2}+B^{3}d)\langle D_{1}\rangle$
and
$\langle T_{2}\rangle=(A^{3}d+3A^{2}B+AB^{2}+2AB^{2}d+B^{3}d^{2})\langle\Theta_{1}\rangle.$
Here we use $B=A^{-1}$ and $d=-A^{2}-A^{-2}$. One can then see that $\langle T_{1}\rangle$ and $\langle T_{2}\rangle$ are sufficiently different to verify that $T_{1}$ and $T_{2}$ are distinct rigid vertex bonded knots. In fact, we specifically calculate that
$\langle T_{1}\rangle=1+1/A^{8}+a/A^{7}+1/A^{6}-1/A^{4}-a/A^{3}-A^{4}-aA^{5}+A^{8}$
and that
$\langle T_{2}\rangle=-1/A^{10}+a/A^{7}-2/A^{4}-1/A^{2}+a/A+aA-A^{4}-aA^{5}+A^{8}.$
This shows explicitly that the two bonded knots are not rigid vertex isotopic.
We note that the highest power of $a$ appearing among all monomials in the expansion of the bonded bracket polynomial of a bonded link diagram equals the total number of bonds. Finally, we note that invariants defined for standard bonded links extend to tight bonded links, and, in some cases, they become simpler to compute due to the absence of crossings between bonds and link arcs.
@@ 5. Bonded Braids
We recall that a geometric braid on $n$ strands is a homeomorphic image of $n$ arcs in the interior of $0,1]\times[0,\epsilon]\times[0,1]$, $\epsilon>0$, such that it is monotonous, that is, there are no local maxima or minima, and the boundary of the image consists in $n$ numbered points in $0,1]\times[0,\epsilon]\times\{0\}$ and $n$ corresponding numbered points in $0,1]\times[0,\epsilon]\times\{1\}$. We study braids through their diagrams in the plane $0,1]\times\{0\}\times[0,1]$, which are also called braids. Formally, a braid is described by its braid word in the braid group $B_{n}$, which has generators $\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1}$ and relations:
@@ 5.1. Bonded braid definition
A bonded braid on $n$ strands is a pair $(\beta,B)$ of a classical braid $\beta$ on $n$ strands, and a set of $k$ disjoint, embedded horizontal simple arcs, called bonds. The boundary points of a bond, called nodes, cannot coincide with endpoints of the braid or with other nodes, and they have local neighborhoods that are 3-valent graphs with the attaching arcs, like $\vdash$, forming together an H-neighborhood, as depicted in Figure 23. A bond joining the $i^{th}$ and $j^{th}$ strands with $i<j$ is denoted $b_{i,j}$ and it threads transversely through the strands in between. By abuse of notation, we denote by $b_{i,j}$ a bond joining the $i^{th}$ and $j^{th}$ strands with any sequence of overpasses or underpasses. A bond $b_{i,i+1}$ joining two consecutive strands $i$ and $i+1$ shall be called elementary bond, and will be denoted by $b_{i}$. If $B=\emptyset$, then $(b,\emptyset)$ is just a classical braid. Moreover, a bonded braid diagram (also referred to as bonded braid) is a regular projection of a bonded braid in the plane $0,1]\times\{0\}\times[0,1]$ with over/under information at every crossing, and such that no crossing is horizontally aligned with a bond. So, $b_{i,j}$ can be encoded unambiguously by a sequence of ji+1j-i+1 o’s and u’s indicating the types of crossings formed between the bond and any strand in between. For an example see left-hand side illustration of Figure 45. A configuration of a bond which passes either over or under all its intermediate strands, such that all crossings are marked ‘over’/‘under’, is a uniform over/under.
@@ 5.2. Bonded braid isotopy
    Figure 47. Interactions between two distant bonds.
@@
  • Furthermore, two bonds $b_{i,j}$ and $b_{k,l}$ commute if $i<k<l<j$, the sequence of crossings of $b_{k,l}$ coincides with a subsequence of the sequence of crossings of $b_{i,j}$ for the strands $k,l$ and the markings in the bigger sequence before and after the subsequence agree, i.e. they are both ‘over’ or both ‘under’. For example, if $b_{k,l}$ forms the sequence $(o,o,u,o)$ with its intermediate strands, the bond $b_{i,j}$ must contain either the subsequence $[o,(o,o,u,o),o]$ or the subsequence $[u,(o,o,u,o),u]$ for the bonds to commute. We call this configuration the matching crossing sequences. For an illustration see Figure 49.
@@
An interaction between a bond and a crossing depends on the relative positions of the bond and the strands forming the crossing. More precisely:
@@
    Figure 51. A crossing commuting with a bond.
@@
  • Figures 54, 55 illustrate the cases where one node of the bond lies on a strand, resp. on two stands, that crosses (resp. cross) another strand. In the instance of Figure 55 there may be other strands threading through the bond, as long as they all lie under the overcrossing strand.
@@
The additional moves depicted in Figures 48, 49 and 53 are essential to define bonded braid equivalence. These moves account for the interactions specific to bonds, ensuring that the topological properties of the bonded braid are preserved. Using these moves, we extend classical braid isotopy to include bonded braid isotopy, leading to the following formal definition.
@@ 5.3. The Bonded Braid Monoid
$\sigma_{1},\ldots,\sigma_{n-1}\quad(\text{classical invertible braid generators})\quad\text{and}\quad b_{i,j}\quad(1\leq i<j\leq n)\quad(\text{bonds}),$
@@
[…]
@@
For reference, the singular braid monoid $SB_{n}$ is generated by the standard braid generators and their inverses $\sigma_{1}^{\pm 1},\ldots,\sigma_{n-1}^{\pm 1}$ together with singular crossing generators $\tau_{1},\ldots,\tau_{n-1}$, subject to the following relations:
[…]
@@
$\begin{array}[]{rcl}b_{2}&=&(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})\\[4.0pt] b_{3}&=&(\sigma_{3}\sigma_{2})^{-1}(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})(\sigma_{3}\sigma_{2})\\[4.0pt] &\vdots&\\[4.0pt] b_{i}&=&(\sigma_{i}\sigma_{i-1})^{-1}\cdots(\sigma_{2}\sigma_{1})^{-1}\,b_{1}\,(\sigma_{2}\sigma_{1})\cdots(\sigma_{i}\sigma_{i-1}),\\[4.0pt]$ $\vdots&\end{array}$
@@
The tight bonded braid monoid $BB_{n}$ admits the following irredundant presentation: it is generated by the classical braid generators and their inverses $\sigma_{1}^{\pm 1},\ldots,\sigma_{n-1}^{\pm 1}$ and a single bond generator $b_{1}$, subject to the relations:
[…]
6. An analogue of the Alexander Theorem for bonded links
As noted in Section 5, one can apply to a bonded braid the usual closure operation and obtain a bonded link (recall Figure 45). In this section we focus on the theory of topological standard bonded links and we present a braiding algorithm for oriented standard bonded links. Namely, we prove the following bonded analogue of the Alexander Theorem for classical links:
@@ 6. An analogue of the Alexander Theorem for bonded links
Every oriented topological standard bonded link can be represented topological vertex isotopically as the closure of a standard resp. tight bonded braid.
@@
Figure 60. Upward oriented arcs are further subdivided to up-arcs, entirely ‘over’ or ‘under’.
@@
Step 2: We next deal with up-arcs in the diagram that contain at least one node of a bond, standard or tight. For this we apply the isotopy moves TVT or RVT as exemplified in Figures 63 and 64, according to whether the two joining arcs are antiparallel or parallel upward arcs. Note that the braiding preparation of a bonded region with two parallel upward arcs can be also achieved with planar isotopy, by performing a 180-degree rotation on the plane. Then the bonds will only lie on down-arcs, so that the braiding algorithm will not affect them.
Figure 63. Braiding preparation for two antiparallel bonded arcs using TVT-moves.
Figure 64. Braiding preparation for two parallel bonded arcs oriented upward using RVT-moves.
@@ 7. Bonded braid equivalences
In this section we formulate and prove two bonded braid equivalences for topological bonded braids: the bonded $L$-equivalence for standard bonded braids and the tight bonded $L$-equivalence for tight bonded braids. For proving a braid equivalence we need to have the diagrams in general position, as described in the proof of the bonded braiding algorithm, to which we now add the triangle condition, whereby any two sliding triangles of up-arcs do not intersect. This is achieved by further subdivision of up-arcs, if needed, and it allows any order in the sequence of the braiding moves. We note here that the interior of a sliding triangle may intersect a bond.
We proceed with introducing the notion of $L$-moves in the bonded setting.In the classical setting, the $L$-moves naturally generalize the stabilization moves in the Markov Theorem for classical braids, since an LL–move is equivalent to adding in a braid a positive or a negative crossing, so that two braids that differ by an $L$-move have isotopic closures. Moreover, as shown in [30], the $L$-moves can also realize conjugation for classical braids. On the other hand, an LL–move can be created by braid isotopy, stabilization and conjugation [30]. The LL–move approach to Markov-type theorems is flexible and powerful for formulating braid equivalences in practically any topological setting. They prove to be particularly useful in settings where the braid analogues do not have an apparent algebraic structure. In [30, 29, 24] $L$-moves and braid equivalence theorems are presented for different knot theories.
@@
An $L$-move on a bonded braid $\beta$, consists in the following operation or its inverse: we cut an arc of $\beta$ open and we pull the upper cutpoint downward and the lower upward, so as to create a new pair of braid strands with corresponding endpoints (on the vertical line of the cutpoint), and such that both strands run entirely over or entirely under the rest of the braid (including the bonds). Pulling the new strands over will give rise to an $L_{o}$-move and pulling under to an $L_{u}$-move. By definition, an $L$-move cannot occur in a bond or at a node. By a general position argument, the new pair of strands does not pass by a crossing or a node. Figure 65 shows $L_{o}$ and $L_{u}$ moves taking place above or below a bond, on its attaching strand.
@@
We note that the closures of two bonded braids that differ by an $L$-move are rigid vertex isotopic. See second illustration of Figure 66, where the closure of the $L$-move contracts to the original arc.
Note 7.3.
@@
Figure 66. A crossing formed after an $L$-move is performed, and pulling the strands of an $L$-move away from the bond.
We consider now the left hand and the right hand side bonded diagrams in Figure 63, which differ only by the one crossing which can be positive or negative. We braid the two bonded diagrams using the braiding moves described in the proof of Theorem 6.1. Examining carefully the resulting bonded braids leads to the following definition.
Definition 7.4.
A bonded $L$-move between two bonded braids resembles an $L$-move with an in-box crossing. More precisely, a bonded $L_{o}$-move (resp. a bonded $L_{u}$-move) consists in the following operation or its inverse: we cut an arc adjacent to a bond node and create with the two cutpoints a crossing of specific type. We then pull the two ends, the upper upward and the lower downward, so as to create a new pair of vertically aligned braid strands, such that both strands run entirely over (resp. entirely under) the rest of the bonded braid (including the bonds). The choice of the crossing is determined by the following property: a vertex slide move with the arc of the crossing not adjacent to the bond node cannot give rise to a classical $L_{o}$-move with a crossing (resp. a classical $L_{u}$-move with a crossing). More precisely, if the two cutpoints lie in the upper right arc of the H-region of a bond, the crossing is positive, while if they lie in the lower right arc of the H-region, the crossing is negative. See Figure 7.4. If the two cutpoints lie in the upper left arc of the H-region of a bond, the crossing is negative, while if they lie in the lower left arc, the crossing is positive.
Furthermore, a tight bonded $L$-move is a bonded $L$-move defined in analogy to a tight $L$-move.
Figure 67. Bonded $L$-moves.
Note 7.5.
As indcated in Figure 67, the closures of two bonded braids that differ by a bonded $L$-move are topologically vertex equivalent.
Definition 7.6.
$L$-moves (resp. tight $L$-moves) together with bonded $L$-moves (resp. tight bonded $L$-moves) and bonded braid isotopy generate an equivalence relation in the set of all bonded braids (resp. tight bonded braids), the bonded $L$-equivalence (resp. tight bonded $L$-equivalence).
@@
Theorem 7.7 (Bonded $L$-equivalence for topological bonded braids).
Two bonded braids upon closure give rise to topologically vertex isotopic oriented standard bonded links if and only if they can be obtained one from the other by a finite sequence of bonded braid isotopy and the following moves:
[…]
Furthermore, two tight bonded braids upon closure give rise to topologically vertex isotopic oriented tight bonded links if and only if they can be obtained one from the other by a finite sequence of tight bonded braid isotopy and the following moves:
[…]
@@
Figure 68. Tight bond commuting.
@@
For the one direction, the closures of two standard/tight bonded braids that differ by the moves of either statement are clearly topologically vertex equivalent. Indeed, the closures of $L$-moves are discussed in Notes 7.2 and 7.3 (recall Figure 66), for the closures of bonded $L$-moves recall Figure 67 and Note 7.5, while bonded commuting is realized via planar isotopy.
For the converse, in order to ensure that the stated moves are sufficient we need to examine any choices made for bringing a bonded diagram to general position and during the braiding algorithm, and show that they result in bonded $L$-equivalent bonded braids. Similarly, that any bonded isotopy moves on a bonded diagram correspond to bonded $L$-equivalent bonded braids. We shall only examine choices involving bonds. All others are proved as in the classical case [30].
@@
Another choice made during the braiding algorithm is when applying a TVT-move, as exemplified in Figure 63, so as to prepare our diagram for braiding. The crossing involved in the move can be positive or negative. One can then easily verify that the resulting braids are bonded $L$-equivalent.
@@
We finally check the vertex slide (VS) moves. The three representing cases of oriented VS moves with the middle arc being an up-arc are illustrated in Figures 73, 74 and 75. The moves are considered local so that all other braiding is done and we can compare the final braids. Figure 73 shows the case of parallel down-arcs at the bonding sites. After the performance of the $L$-moves we see that the move is invisible on the bonded braid level.
@@
The $L$-moves naturally generalize the stabilization moves for classical braids, since an LL–move is equivalent to adding a positive or a negative crossing (recall Note 7.3 and Figure 66). Moreover, as shown in [30], the $L$-moves can also realize conjugation for classical braids and this carries through also to bonded braids. On the other hand, as demonstrated in Figure 66, an LL–move can be created by braid isotopy, stabilization and conjugation. Hence, we may replace moves (1) of Theorem 7.7 and, along with the algebraization of the bonded $L$-moves, we can obtain the analogue of the Markov theorem for bonded braids. This will be discussed in a further paper.